OpenStudy (nerdsarecool):

What is the probability of getting either blue or green on a spinner that is 3/10 green and 1/5 blue? If you were to spin at right. What would be the probability of lading on green or purple?

OpenStudy (nerdsarecool):

Please help i will medal

OpenStudy (anglina4196):

Blue is 2/10 probability wich is 2 out of 10 and green is 3/10 so its 3 out of 10 green

OpenStudy (nerdsarecool):

Ok so 5/10

OpenStudy (anglina4196):

yes

OpenStudy (sshayer):

\[P \left( A \cup B \right)=P \left( A \right)+P \left( B \right)\] \[P \left( A ~ intersection~ B\right)=0\] as A and B are different.(Green and Blue are different.)